Below is the probability distribution for random variable x. What is the probability of at least a score of 2 in this distribution? X 1 2 3 P(x) 0.18 0.42 0.40 a. .82. b. .18. c. .42. d. .60.
Answer #1

Here the probability that at least a score of 2 will be

P(2)+p(3)=0.42+0.4=0.82 option a